The table below gives the CPI basket for 2012. Suppose that 2012 is the reference base period. (5 

marks each)

Item Quantity 

(2012)

Price 

(2012)

Price 

(2013)

Oranges 50 $0.90 $0.75

Bananas 100 $0.50 $0.95

Chicken 200 $2.00 $2.50

Beef 100 $5.00 $4.80

Bread 300 $1.75 $2.00

a) What is the cost of the CPI basket in 2012?

b) What is the cost of the CPI basket in 2013?

c) What is the CPI for 2012?

d) What is the CPI for 2013


1
Expert's answer
2021-08-30T14:35:57-0400

a) Using the given information, the cost of CPI basket in 2012 would be:

= sum of the multiplication of the price of the products in 2012 and the quantity of the products in 2012

"=(50×\$0.90)+(100×\$0.50)+(200×\$2.00)+(100×\$5.00)+(300×\$1.75)\\

=45+50+400+500+525\\

=\$1,520"


b)Similarly, using the given information, the cost of CPI basket in 2013 would be:

= sum of the multiplication of the price of the products in 2013 and the quantity of the products in 2013

"=(50×\$0.75)+(100×\$0.95)+(200×\$2.50)+(100×\$4.80)+(300×\$2.00)\\

=37.5+95+500+480+600\\

=\$1,712.5"


c) here, it is given that the base year is 2012 due to which CPI for 2012 will be 100, such that:

=(Cost of basket in the current year (2012)/cost of basket in base year (2012)]×100

"=(\frac{\$1,520}{\$1,520})×100\\

=100"


d)Similarly, the CPI for 2013 would be:

=(Cost of basket in the current year (2013)/cost of basket in base year (2012)]×100

"=(\frac{\$1,712.5}{\$1,520})×100\\

=112.66"